# Learning Low-Dimensional Embeddings for Black-Box Optimization

### v1

@@ Abstract
When gradient-based methods are impractical, black-box optimization (BBO) provides a valuable alternative. However, BBO often struggles with high-dimensional problems and limited trial budgets. In this work, we propose a novel approach based on meta-learning to pre-compute a reduced-dimensional manifold where optimal points lie for a specific class of optimization problems. When optimizing a new problem instance sampled from the class, black-box optimization is carried out in the reduced-dimensional space, effectively reducing the effort required for finding near-optimal solutions.
@@ 1 Introduction
Black-box optimization (BBO) aims to find the solution of an optimization problem where the objective function is unknown or lacks an explicit mathematical formulation. Instead, the function can only be evaluated through queries, such as physical experiments or simulations through complex computational models. However, in many real-world scenarios, these evaluations are expensive, noisy, or time-consuming. Therefore, a key goal of black-box optimization is to find near-optimal solutions while limiting the number of costly function evaluations. Due to these challenges, BBO relies on sample-efficient strategies that select query points to balance exploration (searching unexplored regions) and exploitation (refining promising solutions). Various techniques exist for tackling these problems, including Kriging-based methods [6], Bayesian Optimization (BO) [12], Mesh adaptive direct search [1], and recently proposed algorithms such as Set-membership approaches [10] and global optimization using inverse distance weighting and surrogate radial basis functions (GLIS) [2].
@@
The primary limitation of the BBO algorithms mentioned above is their scalability with respect of the number of decision variables, restricting their applicability to small- to medium-size optimization problems. Black-box optimization in high-dimensional spaces still remains an open challenge, as the number of function evaluations required to cover the domain of decision variables increases exponentially with the number of dimensions. In recent years, there have been a few works to mitigate this curse of dimensionality issue, aiming at developing methods that enhance the scalability of BBO in high dimensions by exploiting different assumptions about the structure of the underlying objective function. The seminal work Random EMbedding BO (REMBO) [14] laid the foundation for addressing Bayesian Optimization in high-dimensional spaces. REMBO leverages the low effective dimensionality of the objective function and employs a random projection matrix to map the decision variables into a lower-dimensional subspace, where execution of BO is feasible. However, it has been shown that this approach may perform poorly even on some synthetic problems with a true low-dimensional linear subspace due to distortions introduced by the random linear projections [8]. Alternatively, the methods in [7] and [9] assume an additive structure of the objective function and decompose it into a sum of low-dimensional components, each depending on a small subset of variables. In [15], data-independent rules are explored to identify effective additive decompositions.
@@
By leveraging the meta-learning paradigm [11], our paper focuses on reducing the dimension of the decision space where BBO is performed. Meta-learning, or “learning to learn”, involves acquiring knowledge from a class of related tasks to improve the learning process for new tasks within the same class. Following this principle, the key idea behind our approach is to learn a low-dimensional manifold where the optimal solutions of a class of similar optimization problems lie. To identify this manifold, we learn an embedding that maps the optimal decision variables of a set of problems within the considered class into a lower-dimensional latent space. This embedding is obtained using an autoencoder (AE) neural network, trained to minimize the reconstruction error between the optimal solutions in the original higher-dimensional space and the decoder output. The latent space of the trained AE thus serves as the desired low-dimensional representation. The problem addressed in the paper is further motivated by considering as a problem-class a multiparametric programming problem [4]. Under suitable assumptions, the optimizer can be written explicitly as a function of the parameter vector. This defines a manifold where all possible optimizers lie.
@@
An advantage of our approach is that the optimal solutions in the original input domain can be precomputed offline using, for example, simulators with varying parameters and computationally expensive optimization algorithms, e.g., evolutionary algorithms. This precomputed knowledge is then leveraged to solve new optimization problems by performing black-box optimization directly in the learned embedded space rather than the original higher-dimensional space. For instance, in tuning of Model Predictive Control (MPC) hyperparameters, one can first compute optimal hyperparameters using simulators of a class of parameterized systems. Then, when applying the controller to a real system, calibration is performed efficiently through experiments using BBO algorithms such as Bayesian Optimization or GLIS, searching over the precomputed embedded space rather than the original higher-dimensional domain. Although our results can be applied, in general, to any BBO method, in this paper we primarily focus on the GLIS algorithm.
@@ 2 Problem description
$J(\theta)=\min_{x}f(x;\theta)\qquad\text{s.t.}\quad x\in\mathcal{X}$ (1)
where, for a given parameter $\theta$, $f(x;\theta):\mathbb{R}^{n_{x}}\to\mathbb{R}$ is a function of the optimization vector $x\in\mathcal{X}$. The parameter $\theta$ represents a set of conditions that affect the optimization problem. These conditions may correspond to simulation settings, system configurations, or operational scenarios in real-world applications, which in turn influence the structure of the function $f(x;\theta)$. While $\theta$ is here represented as a vector in $\mathbb{R}^{n_{\theta}}$, it should be understood as a description of the variations in the environment or in the system that modify the cost function, rather than merely as a vector of numbers.
We assume that an analytical expression of $f(x;\theta)$ is not available, but the function can be evaluated at any query input $x$ and for any parameter $\theta$. This may be the case, for instance, when $\theta$ describes different configurations of the environment and $x$ is a design variable to be computed in order to optimize a metric $f$, which in turn can be evaluated only through execution of real experiments. The set $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ represents known constraints on $x$, such as box or nonlinear constraints. The parameter $\theta$ is assumed to have a given prior distribution $p(\theta)$ from which its values can be sampled.
Let us now consider a new optimization variable $z\in\mathbb{R}^{n_{z}}$, with $n_{z}<n_{x}$. The problem addressed in this work is to learn a map $\mathrm{D}:\mathbb{R}^{n_{z}}\to\mathbb{R}^{n_{x}}$ such that the optimization of $f$ can be carried out effectively over the lower-dimensional variable $z$. Problem (1) is then transformed to:
$J(\theta)=\min_{z}f(\mathrm{D}(z);\theta)\qquad\text{s.t.}\quad z\in\mathcal{Z},$ (2)
where the embedding satisfies the constraints $\mathcal{Z}=\{z\in\mathbb{R}^{n_{z}}:\mathrm{D}(z)\in\mathcal{X}\}$. A good map $D$ should be such that the solution $x^{\star}(\theta)=\arg\min_{x\in\mathcal{X}}f(x;\theta)$ of the original problem (1) can be recovered from the solution $z^{\star}(\theta)$ of the reduced dimensional problem (2) as $x^{\star}(\theta)\approx\mathrm{D}(z^{\star}(\theta))$.
@@ 3.1 Meta-dataset creation
We generate a set of $N$ parameters $\Theta:=\{\theta^{(i)}\}_{i=1}^{N}$ by sampling from the prior parameter distribution, i.e., $\theta^{(i)}\sim p(\theta)$. For each $\theta^{(i)}$, we solve (1) with any sample-based optimization algorithm (such evolutionary learning, MCMC with simulated annealing, GLIS) and we gather a set $\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{K}$of $K$ near-optimal vectors $x^{(i)}_{k}$, and the corresponding function values $f^{(i)}_{k}=f(x^{(i)}_{k},\theta^{(i)})$, with $x^{(i)}_{l}\neq x^{(i)}_{m},\forall l\neq m$, $l,m=1,\ldots,K$. For a given $\theta^{(i)}$, we refer to as near-optimal vectors as the best $K$ candidate points explored during the iterations of the numerical algorithms used to solve (1). We denote the meta-dataset as
@@
3.2 Learning latent representation with auto encoders
An autoencoder (AE) is a type of neural network consisting of an encoder and a decoder, designed to learn a compressed representation of data [5]. It is trained to map an input vector $x\in\mathbb{R}^{n_{x}}$ to itself, while passing through a bottleneck layer of size $n_{z}<n_{x}$. The idea behind AEs is that the network learns a latent description of the input in a lower-dimensional latent space $\mathbb{R}^{n_{z}}$.
Let us consider an autoencoder $\mathcal{AE}_{\phi}:\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ with tunable parameters $\phi$, which is composed of an encoder $\mathrm{E}_{\phi_{e}}:\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{z}}$ mapping $x$ to a lower dimensional latent space vector $z$ and a decoder $\mathrm{D}_{\phi_{d}}:\mathbb{R}^{n_{z}}\to\mathbb{R}^{n_{x}}$ mapping the latent vector $z$ to reconstruct the original vector $x$. Given the meta-dataset $\mathcal{D}$ in (3), we train $\mathcal{AE}_{\phi}$ by minimizing the following weighted reconstruction error over the parameters $\phi=\{\phi_{e},\phi_{d}\}$,
$\mathcal{L}(\phi)=\min_{\phi}\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}w^{(i)}_{k}\|x^{(i)}_{k}-\mathcal{AE}_{\phi}(x^{(i)}_{k})\|^{2}_{2}.$ (4)
Given a hyperparameter $\alpha\in[0,1)$, the weights $w^{(i)}_{k}$ are defined as
$w^{(i)}_{k}=\alpha^{\mathrm{argsort}(f_{k}^{(i)},\;\{f_{h}^{(i)}\}_{h=1}^{K})},$ (5)
with the function $\mathrm{argsort}(\cdot,\cdot)$ returning the index of its first argument in sorted order, from smallest to largest, with respect to the elements in the second argument. This weighting gives relatively more importance to the best candidate points $x_{k}^{(i)}$.
After completing the training, black-box optimization is performed over $z$ in a lower-dimensional space. The decoder then maps the latent variable $z$ back to the original high-dimensional space, where the function $f(x;\theta)$ is evaluated. In the following paragraph, we provide more details on the application of the proposed approach for the implementation of GLIS in a low-dimensional space. Nevertheless, the same ideas can also be applied to other BBO algorithms, such as Bayesian Optimization.
@@ 3.2 Learning latent representation with auto encoders
In order to ensure that the autoencoder output lies within the feasible set $\mathcal{X}$, penalty terms can be added to the loss function in (4). Alternatively, the last layer of the decoder can be designed to enforce constraint satisfaction. For instance, in the case of box constraints, sigmoid activation functions can be used in the last layer of the decoder, and their outputs can then be rescaled based on the specified bounds.
@@
[…]
@@ 3.3 Meta-GLIS: GLIS over a low-dimensional space
In this section, we describe how to apply the proposed methodology to implement Meta-GLIS—a lightweight version of GLIS based on meta learning that operates in a low-dimensional latent space. First, we recall that the encoder $\mathrm{E}_{\phi_{e}}$ and decoder $\mathrm{D}_{\phi_{d}}$ are pretrained (e.g., in a simulation environment) using the meta-dataset $\mathcal{D}$ defined in (3). Given a query parameter $\bar{\theta}$ (which represents, for instance, the conditions of a real-world problem, rather than a simulated one), the goal is to solve the original optimization problem (1) for $\theta=\bar{\theta}$. Outside the simulation environment, evaluating the objective function $f(x,\bar{\theta})$ is assumed to be costly. Therefore, Meta-GLIS is expected to provide a near-optimal solution within a limited number of function evaluations.
[…]
@@
where $h(z)$ is an exploration-promoting function, and $\delta>0$ is a tuning hyperparameter that balances exploitation and exploration in the acquisition function. According to GLIS [2, Section 4], the exploration term is an Inverse Distance Weighting (IDW) function, defined as:
$h(z)=\begin{cases}0,&\text{if }z\in\mathcal{D}_{M},\\ \frac{2}{\pi}\tan^{-1}\left(\frac{1}{\sum_{i=1}^{M}r_{i}(z)}\right),&\text{otherwise},\end{cases}$ (7)
where $r_{i}(z)=\frac{1}{\|z-z_{i}\|^{2}}$, with $z_{i}\in\mathcal{D}_{M}$. The acquisition function is then minimized to determine the next sampling point $z_{M+1}$ and the corresponding decoded value $x_{M+1}=\mathrm{D}_{\phi_{d}}(z_{M+1})$, at which the objective function is evaluated (Steps 4.3 and 4.4). The process continues until the maximum number of iterations is reached. The best solution is then returned.
@@
Input: number of initial samples $M_{\text{init}}$; maximum number of function evaluations $M_{\max}\geq M_{\text{init}}$; pre-trained encoder $\mathrm{E}_{\phi_{e}}$ and decoder $\mathrm{D}_{\phi_{d}}$; query parameter $\bar{\theta}$.
  1. 1.
    Set $M\leftarrow M_{\text{init}}$;
  2. 2.
    Generate $M$ random initial samples $Z=[z_{1},\dots,z_{M}]$ using Latin hypercube sampling from $\mathcal{Z}$;
  3. 3.
    Create the dataset $\mathcal{D}_{M}=\{z_{j},f(\mathrm{D}_{\phi_{d}}(z_{j};\bar{\theta}))\}_{j=1}^{M}$;
  4. 4.
    while $M<M_{\max}$
    1. 0..1.
      Compute a surrogate $\hat{f}(z)$ of $f(\mathrm{D}_{\phi_{d}}(z);\bar{\theta})$ from $\mathcal{D}_{M}$;
    2. 0..2.
      Compute the acquisition function $a(z)$ (eq. (6));
    3. 0..3.
      Compute next input sample $z_{M+1}=\arg\min\limits_{z\in\mathcal{Z}}a(z)$;
    4. 0..4.
      Evaluate the function $f(\mathrm{D}_{\phi_{d}}(z_{M+1});\bar{\theta})$;
    5. 0..5.
@@
      $\mathcal{D}_{M+1}=\mathcal{D}_{M}\cup\{z_{M+1},f(\mathrm{D}_{\phi_{d}}(z_{M+1});\bar{\theta})\}$
    6. 0..6.
@@
  5. 5.
    end while
  6. 6.
    Compute the best solution:
    $z^{*}=\arg\min_{z\in\mathcal{D}_{M}}f(\mathrm{D}_{\phi_{d}}(z);\bar{\theta})$
    $x^{*}=\mathrm{D}_{\phi_{d}}(z^{*}),\ \ \ f^{*}=f(x^{*};\bar{\theta})$

### v2

@@ Abstract
Black-box optimization (BBO) is a valuable alternative to gradient-based methods when computing derivatives of the objective function is not feasible. However, BBO often struggles with high-dimensional decision spaces and limited evaluation budgets. In this work, we propose a meta-learning approach that pre-computes a reduced-dimensional manifold capturing the optimal solutions of a class of optimization problems. For new problem instances drawn from the same class, BBO is then performed in the latent space, substantially reducing the computational effort required to reach near-optimal solutions. We further establish probabilistic upper bounds on the performance gap between solving the original problem and its reduced-dimensional counterpart. The effectiveness of the method is demonstrated both on benchmark optimization problems and on the automatic tuning of model predictive controllers.
@@ 1 Introduction
Black-box optimization (BBO) aims to find the solution of an optimization problem where the objective function is unknown or lacks an explicit mathematical formulation. Instead, the function can only be evaluated through queries, such as physical experiments or simulations through complex computational models. However, in many real-world scenarios, these evaluations are expensive, noisy, or time-consuming. Therefore, a key goal of black-box optimization is to find near-optimal solutions while limiting the number of costly function evaluations. Due to these challenges, BBO relies on sample-efficient strategies that select query points to balance exploration (searching unexplored regions) and exploitation (refining promising solutions). Various techniques exist for tackling these problems, including Kriging-based methods [7], Bayesian Optimization (BO) [17], Mesh adaptive direct search [1], and recently proposed algorithms such as Set-Membership (SM) approaches [15] and global optimization using inverse distance weighting and surrogate radial basis functions (GLIS) [2].
@@
The primary limitation of the BBO algorithms mentioned above is their scalability with respect to the number of decision variables, restricting their applicability to small- to medium-size optimization problems. Black-box optimization in high-dimensional spaces still remains an open challenge, as the number of function evaluations required to cover the domain of decision variables increases exponentially with the number of dimensions. In recent years, there have been a few works to mitigate this curse of dimensionality issue, aiming at developing methods that enhance the scalability of BBO in high dimensions by exploiting different assumptions about the structure of the underlying objective function. The seminal work Random EMbedding BO (REMBO) [19] laid the foundation for addressing Bayesian Optimization in high-dimensional spaces. REMBO leverages the low effective dimensionality of the objective function and employs a random projection matrix to map the decision variables into a lower-dimensional subspace, where execution of BO is feasible. However, it has been shown that this approach may perform poorly even on some synthetic problems with a true low-dimensional linear subspace due to distortions introduced by the random linear projections [9]. Alternatively, the methods in [8] and [12] assume an additive structure of the objective function and decompose it into a sum of low-dimensional components, each depending on a small subset of variables. In [20], data-independent rules are explored to identify effective additive decompositions. In [14], Bayesian optimization is implemented iteratively by optimizing a subset of parameters while keeping the others fixed. The methodology is applied to tune the PID gains of a torque-controlled, 7-degree-of-freedom robot manipulator.
@@
By leveraging the meta-learning paradigm [16], our paper focuses on reducing the dimension of the decision space where BBO is performed. Meta-learning, or “learning to learn”, involves acquiring knowledge from a class of related tasks to improve the learning process for new tasks within the same class. Following this principle, the key idea behind our approach is to learn a low-dimensional manifold where the optimal solutions of a class of similar optimization problems lie. To identify this manifold, we learn an embedding that maps the optimal decision variables of a set of problems within the considered class into a lower-dimensional latent space. This embedding is obtained using an autoencoder (AE) neural network, trained to minimize the reconstruction error between the optimal solutions in the original higher-dimensional space and the decoder output. The latent space of the trained AE thus serves as the desired low-dimensional representation.
@@
An advantage of our approach is that the optimal solutions in the original input domain can be precomputed offline using, for example, simulators with varying parameters and computationally expensive optimization algorithms, e.g., evolutionary algorithms. This precomputed knowledge is then leveraged to solve new optimization problems by performing black-box optimization directly in the learned embedded space rather than the original higher-dimensional space. For instance, in tuning of Model Predictive Control (MPC) hyperparameters, one can first compute optimal hyperparameters using simulators of a class of parameterized systems. Then, when applying the controller to a real system, calibration is performed efficiently through experiments using BBO algorithms such as Bayesian Optimization or GLIS, searching over the precomputed embedded space rather than the original higher-dimensional domain. Although our methodology can, in principle, be applied in conjunction with any BBO method, in this paper we primarily focus on the GLIS algorithm. The problem addressed in the paper is further motivated by considering a multiparametric programming problem [4] as a problem-class. Under suitable assumptions, the optimizer can be written explicitly as a function of the parameter vector, thereby defining a manifold where all the optimizers lie.
As a further contribution, in this work we provide a theoretical result that establishes a probabilistic, distribution-free bound on the performance loss incurred when solving an optimization problem in the latent space instead of the original domain. The result does not depend on the specific architecture used to construct the embedding, and can in principle be applied to neural autoencoders, random projections, or other dimensionality-reduction techniques. Furthermore, the bound does not compare the (generally unattainable) global optima of the two problems, but rather the solutions actually returned by the numerical algorithms employed in each space. This bound is independent of the particular numerical optimization algorithms employed in the two spaces. For example, optimization in the original domain could be carried out with gradient-based methods or Particle Swarm Optimization, while the reduced problem may be solved using GLIS or other black-box algorithms.
The rest of the paper is organized as follows. In Section 2, we introduce the class of black-box optimization problems considered in this work and formalize their reduced-dimensional counterparts. Section 3.2 presents the proposed methodology for learning low-dimensional embeddings via autoencoders trained on meta-datasets of near-optimal solutions. Section 3.3 describes the Meta-GLIS algorithm, which applies the GLIS method in the learned latent space. In Section 4, we provide theoretical results quantifying the sub-optimality gap between optimization in the latent and original domains. Section 5 reports numerical results obtained on a nonconvex optimization benchmark and for the automated calibration of MPC hyperparameters. Finally, Section 6 draws conclusions and discusses directions for future work.
@@ 2 Problem description
$f^{\star}_{x}(\theta)=\min_{x}f(x;\theta)\qquad\text{s.t.}\quad x\in\mathcal{X}$ (1)
where, for a given parameter $\theta\in\mathbb{R}^{n_{\theta}}$, $f(x;\theta):\mathbb{R}^{n_{x}}\to\mathbb{R}$ is a function of the optimization variable $x\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$. The parameter $\theta$ represents a set of conditions that affect the optimization problem. These conditions may correspond to simulation settings, system configurations, or operational scenarios in real-world applications, which in turn influence the structure of the function $f(x;\theta)$.
We assume that an analytical expression of $f(x;\theta)$ is not available, but the function can be evaluated at any feasible query input $x\in\mathcal{X}$ and for any parameter $\theta$. This may be the case, for instance, when $\theta$ describes different configurations of the environment and $x$ is a design variable to be computed in order to optimize a metric $f$, which in turn can be evaluated only through the execution of a real experiment or high-fidelity simulation. The set $\mathcal{X}$ represents known constraints on $x$. For the theoretical analysis of our approach, we will denote by $p(\theta)$ the distribution from which the values of the parameter $\theta$ are sampled (e.g., a uniform distribution over a hyperbox of parameters). For the practical implementation, as clarified in Section 3, we only assume access to a finite set of samples drawn from $p(\theta)$.
Problem Statement
Given the class of optimization problems (1) in $n_{x}$ variables and a desired reduced dimension $n_{z}<n_{x}$, learn a map $\mathrm{D}:\mathbb{R}^{n_{z}}\to\mathbb{R}^{n_{x}}$ such that the optimization of $f$ can be carried out effectively with respect to the lower-dimensional embedding $z\in\mathbb{R}^{n_{z}}$:
$z^{\star}(\theta)\in\arg\min_{z\in\mathcal{Z}}f(\mathrm{D}(z);\theta),\\$ (2)
with
$f(\mathrm{D}(z^{\star}(\theta));\theta)\approx f^{\star}_{x}(\theta),\quad\forall\theta\in\mathbb{R}^{n_{\theta}},$
where the feasible set $\mathcal{Z}=\{z\in\mathbb{R}^{n_{z}}:\mathrm{D}(z)\in\mathcal{X}\}$.
In other words, our goal is to find a map $\mathrm{D}$ such that we can retrieve an approximation $\mathrm{D}(z^{\star}(\theta))$ of an optimizer $x^{\star}(\theta)$ of the original problem (1) from the solution $z^{\star}(\theta)$ of the simpler problem (2).
@@ 3.1 Meta-dataset creation
We generate a set of $N$ parameters $\{\theta^{(i)}\}_{i=1}^{N}$ by sampling from the parameter distribution, i.e., $\theta^{(i)}\sim p(\theta)$. For each $\theta^{(i)}$, we solve (1) by using any sample-based optimization algorithm (such evolutionary learning, MCMC with simulated annealing, Bayesian Optimization, GLIS) and we gather a set $\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{K}$of $K$ near-optimal vectors $x^{(i)}_{k}$, and the corresponding function values $f^{(i)}_{k}=f(x^{(i)}_{k},\theta^{(i)})$, with $x^{(i)}_{l}\neq x^{(i)}_{m},\forall l\neq m$, $l,m=1,\ldots,K$. For a given $\theta^{(i)}$, we refer to as near-optimal vectors as the best $K$ candidate points explored during the iterations of the numerical algorithms used to solve (1). We denote the meta-dataset as
@@
3.2 Learning a latent representation via autoencoders
An autoencoder (AE) [6] is trained to map an input vector $x\in\mathbb{R}^{n_{x}}$ to itself, while passing through a bottleneck layer of size $n_{z}<n_{x}$. Let us consider an autoencoder $\mathcal{AE}_{\phi}:\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ with tunable parameters $\phi$, which is composed of an encoder $\mathrm{E}_{\phi_{e}}:\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{z}}$ mapping $x$ to a lower dimensional latent space vector $z$ and a decoder $\mathrm{D}_{\phi_{d}}:\mathbb{R}^{n_{z}}\to\mathbb{R}^{n_{x}}$ mapping the latent vector $z$ to reconstruct the original vector $x$. Given the meta-dataset $\mathcal{D}$ in (3), we train the autoencoder $\mathcal{AE}_{\phi}$ by minimizing the following weighted reconstruction error over the parameters $\phi=\{\phi_{e},\phi_{d}\}$:
$\mathcal{L}(\phi)=\min_{\phi}\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}w^{(i)}_{k}\|x^{(i)}_{k}-\mathcal{AE}_{\phi}(x^{(i)}_{k})\|^{2}_{2},$ (4)
where the weights $w^{(i)}_{k}$ are chosen to give relatively more importance to the best candidate points $x_{k}^{(i)}$. In particular, given a hyperparameter $\lambda\in[0,1)$, we define the weights as follows:
$(j_{1}^{(i)},\ldots,j^{(i)}_{K})$ $=\mathrm{argsort}\left(\{f_{k}^{(i)}\}_{k=1}^{K}\right),$
$w^{(i)}_{k}$ $=\lambda^{j_{k}^{(i)}},\quad k=1,\ldots,K,$
where the function $\mathrm{argsort}(\cdot)$ sorts the argument values in ascending order.
After training the autoencoder, black-box optimization is performed over $z$ in a lower-dimensional space. The decoder $\mathrm{D}_{\phi_{d}}$ then maps the latent variable $z$ back to the original high-dimensional space, where the function $f(x;\theta)$ is evaluated. In the following paragraph, we provide more details on the application of the proposed approach for the implementation of GLIS in a low-dimensional space. Nevertheless, the same ideas can also be applied to other BBO algorithms, such as Bayesian Optimization.
@@ 3.2 Learning a latent representation via autoencoders
In order to ensure that the output of the autoencoder $\mathcal{AE}_{\phi}$ lies within the feasible set $\mathcal{X}$, penalty terms can be added to the loss function in (4). Alternatively, the last layer of the decoder $\mathrm{D}_{\phi_{d}}$ can be designed to enforce constraint satisfaction. For instance, in the case of box constraints, scaled and shifted sigmoid activation functions can be used in the last layer of the decoder to enforce the specified bounds.
@@
Similarly to the discussion in Remark 1, bounds on the latent variables can also be enforced. In particular, a sigmoid activation function is used in the last layer of the encoder to rescale the latent variable $z$ such that it belongs to the box $\mathcal{Z}=[0,\ 1]^{n_{z}}$. This restricts the search space over the latent variable domain, as discussed in the next section.
@@ 3.3 Meta-GLIS: GLIS over a low-dimensional space
In this section, we describe how to apply the proposed methodology to implement Meta-GLIS, a lightweight version of GLIS [2] based on meta learning that operates in a low-dimensional latent space. First, we recall that the encoder $\mathrm{E}_{\phi_{e}}$ and decoder $\mathrm{D}_{\phi_{d}}$ are pretrained (e.g., in a simulation environment) using the meta-dataset $\mathcal{D}$ defined in (3). Given a query parameter $\bar{\theta}$ (which represents, for instance, the conditions of a real-world problem, rather than a simulated one), the goal is to solve the original optimization problem (1) for $\theta=\bar{\theta}$. Outside the simulation environment, evaluating the objective function $f(x,\bar{\theta})$ is assumed to be costly. Therefore, Meta-GLIS is expected to provide a near-optimal solution within a limited number of function evaluations.
The Meta-GLIS algorithm, summarized in Algorithm 1, follows a standard surrogate-based black-box optimization strategy. First, an initial set of samples of $M_{\text{init}}$ samples is drawn from the latent space domain $\mathcal{Z}=[0,\ 1]^{n_{z}}$ using Latin Hypercube Sampling (LHS) [11] (Step 2). The function $f(\mathrm{D}_{\phi_{d}}(z);\bar{\theta})$ is then evaluated at these initial points, forming the initial dataset $\mathcal{D}_{M}$ (Step 3). In the iterative optimization loop (Step 4), a surrogate model $\hat{f}:\mathcal{Z}\rightarrow\mathbb{R}$ is constructed to approximate the objective function based on the dataset $\mathcal{D}_{M}=\{z_{j},f(\mathrm{D}_{\phi_{d}}(z_{j});\bar{\theta})\}_{j=1}^{M}$ (Step 4.1). As in the original formulation of GLIS, a linear combination of Radial Basis Functions (RBFs) [5] is used to create the surrogate function. Specifically,
$\hat{f}(z)=\sum_{k=1}^{M}\beta_{j}\phi(\mu d(z,z_{k})),$ (5)
where $\phi:\mathbb{R}\to\mathbb{R}$ is an RBF, $d:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ is the squared Euclidean distance $d(z,z_{k})=\|z-z_{k}\|_{2}^{2}$, $\mu>0$ is a scalar hyper-parameter defining the shape of the RBF. The unknown coefficients $\beta=[\beta_{1}\ \ldots\ \beta_{M}]^{T}$ are computed by fitting the surrogate $\hat{f}(z)$ to the dataset $\mathcal{D}_{M}$ through the minimization of the regularized squared error:
$\sum_{j=1}^{M}\Big\|f(\mathrm{D}_{\phi_{d}}(z_{j};\bar{\theta}))-\underset{\hat{f}(z_{j})}{\underbrace{\sum_{k=1}^{M}\beta_{j}\phi(\mu d(z_{j},z_{k}))}}\Big\|^{2}+\gamma\left\|\beta_{j}\right\|^{2},$ (6)
where the quadratic regularization term (weighted by the hyperparameter $\gamma>0$) is added to guarantee strict convexity of (6). Some RBFs commonly used are $\phi(\mu d)=\frac{1}{1+(\mu d)^{2}}$ (inverse quadratic) and $\phi(\mu d)=e^{-(\mu d)^{2}}$ (squared exponential kernel). Both the hyperparameter $\mu$ and $\gamma$ are tuned through cross-validation.
The acquisition function $a(z)$ is then constructed (Step 4.2) as:
@@
where $h(z)$ is an exploration-promoting function, and $\delta>0$ is a tuning hyperparameter that balances exploitation and exploration in the acquisition function. According to GLIS [2, Section 4], the exploration term is defined by the following Inverse Distance Weighting (IDW) function
$h(z)=\begin{cases}0&\text{if }z\in\mathcal{D}_{M},\\ \frac{2}{\pi}\tan^{-1}\left(\frac{1}{\sum_{i=1}^{M}\frac{1}{\|z-z_{i}\|^{2}}}\right)&\text{otherwise}.\end{cases}$ (8)
The acquisition function (7) is then minimized to determine the next sampling point $z_{M+1}$ and the corresponding decoded value $x_{M+1}=\mathrm{D}_{\phi_{d}}(z_{M+1})$, at which the objective function is evaluated (Steps 4.3 and 4.4). The process continues until the maximum number of iterations is reached. The best solution is then returned.
@@
Input: number $M_{\text{init}}$ of initial samples; maximum number $M_{\max}\geq M_{\text{init}}$ of function evaluations; pre-trained encoder $\mathrm{E}_{\phi_{e}}$ and decoder $\mathrm{D}_{\phi_{d}}$; query parameter $\bar{\theta}$.
1:Set $M\leftarrow M_{\text{init}}$;
2:Generate $M$ random initial samples $Z=[z_{1},\dots,z_{M}]$ using LHS from $\mathcal{Z}$;
3:Create dataset $\mathcal{D}_{M}\leftarrow\{(z_{j},f(\mathrm{D}_{\phi_{d}}(z_{j};\bar{\theta})))\}_{j=1}^{M}$;
4:while $M<M_{\max}$ do:
  1. 4.1
    Compute surrogate $\hat{f}(z)$ of $f(\mathrm{D}_{\phi_{d}}(z);\bar{\theta})$ from $\mathcal{D}_{M}$;
  2. 4.2
    Compute acquisition function $a(z)$ (see Eq. (7));
  3. 4.3
    Compute next input $z_{M+1}=\arg\min\limits_{z\in\mathcal{Z}}a(z)$;
  4. 4.4
    Evaluate $f(\mathrm{D}_{\phi_{d}}(z_{M+1});\bar{\theta})$;
  5. 4.5
@@
    $\mathcal{D}_{M+1}\leftarrow\mathcal{D}_{M}\cup\{(z_{M+1},f(\mathrm{D}_{\phi_{d}}(z_{M+1});\bar{\theta}))\}$
  6. 4.6
@@
5:end while;
6:Compute best solution:
$z^{*}$ $=\arg\min_{z\in\mathcal{D}_{M}}f(\mathrm{D}_{\phi_{d}}(z);\bar{\theta}),$
$x^{*}$ $\leftarrow\mathrm{D}_{\phi_{d}}(z^{*}),\ f^{*}\leftarrow f(x^{*};\bar{\theta}).$